# Design of a CsI(Tl) Calorimeter for Muonium-to-Antimuonium Conversion Experiment

### v2

@@ Abstract
The Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for charged lepton flavor violation and increase the sensitivity by three orders of magnitude compared to the PSI experiment in the 1990s. A clear signature of this conversion is the positron produced from antimuonium decay. This paper presents a near-$4\pi$-coverage calorimeter designed for MACE, which can provide an energy resolution of 9% at 511 keV, and a signal efficiency of 55.5% for annihilation $\gamma$-ray events at maximum. Detailed Monte-Carlo simulations using MACE offline software based on Geant4 are performed for geometry optimization, coincidence system design, background estimation, and benchmark detector validation.
@@ I Introduction
Muonium (M) is an exotic atom consisting of a bound state of $\mu^{+}e^{-}$, while its antiatom, antimuonium ($\bar{\text{M}}$), is composed of $\mu^{-}e^{+}$. From the perspective of the Standard Model Effective Field Theory (SMEFT), the spontaneous M-to-$\bar{\text{M}}$ conversion ($\mu^{+}e^{-}\rightarrow\mu^{-}e^{+}$), which violates lepton flavor by two units instead of one, can probe distinct effective operators and thereby constrain a different parameter space [8, 9, 10]. As a result, the key innovation in investigating M-to-$\bar{\text{M}}$ conversion is providing additional constraints on specific new physics models, distinct from other CLFV processes [11, 12, 13]. The most recent measured upper limit ($8.3\times 10^{-11}$ at 90% C.L.) for the conversion probability in a 0.1 T magnetic field was reported by the MACS experiment at PSI in 1998 [14, 15]. Another experiment focused on M-to-$\bar{\text{M}}$ conversion was proposed and conceptually designed at PRISM in 2003, as reported in [16], but did not progress to further studies. There is also a proposal at J-PARC in Japan to search for this process using a novel scheme [17]. Surprisingly, the limit of M-to-$\bar{\text{M}}$ conversion has remained unchallenged for two decades. There is an urgent need for a new-generation experiment.
In light of significant advancements in muon beamlines and particle detector techniques nowadays, the Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for M-to-$\bar{\text{M}}$ conversion and has the potential to improve the upper limit by three orders of magnitude [18]. With the construction of the China advanced NUclear physics research Facility (CNUF) [19], MACE is expected to be the first muon-CLFV experiment in China, providing extra perspectives compared to other concerned experiments. The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Microchannel Plate (MCP) [20, 21], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence. A schematic of the MACE apparatus is shown in Fig. 1.
@@
In the MACE concept, surface muons with a momentum of 26.3 MeV$/c$ are stopped at a production target where M is formed and then possibly diffuses into the vacuum and converts into $\bar{\text{M}}$. The decay products of $\bar{\text{M}}$ include an electron with Michel energy (up to 52.8 MeV) and a positron with atomic-shell energy ($\sim 13.5$ eV). The Michel electron will be reconstructed by MMS which is expected to achieve $10^{-4}$ level suppression of background in the ECAL end through preliminary simulation. On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the solenoid with transverse-position conservation to MCP, where the positron annihilates into two back-to-back 511 keV $\gamma$-rays. The cylindrical magnet with a field strength of 0.1 T outside ECAL is used to constrain the positron beam spot after transportation. As a rare-process searching experiment, MACE requires high sensitivity for the annihilation $\gamma$-rays to discriminate the signal from backgrounds and retain as much event information as possible. Thus, an ECAL with good energy resolution and high efficiency is essential for the design of MACE.
@@
The main purpose of this study is to propose a design for the MACE $4\pi$ calorimeter using the Monte-Carlo simulation method. In this work, the geometry of ECAL is generated and optimized using the MACE offline software based on the Geant4 toolkit [35, 36]. Subsequently, the energy resolution, signal efficiency, and background level of ECAL are estimated, providing a reference for the MACE conceptual design. This paper is structured as follows Section I introduces the MACE calorimeter design and geometry optimization. Section II presents the simulation results for both signal and background. Section III describes the experiment setups of a benchmark detector to validate the energy resolution study in the MACE calorimeter design. The paper concludes with a discussion of the conclusions and perspectives for future work in Section IV.
@@ II.1 Overview
To meet the requirements of enhanced sensitivity (e.g., excellent energy resolution, high detection efficiency, and good geometrical acceptance), we propose a near-$4\pi$-coverage polyhedral calorimeter design. In this section, we present a design scheme for ECAL, including details on scintillation materials and photosensors. This proposed design will serve as the basis for the simulation work discussed in the following sections.
@@
Calorimeters with inorganic scintillation crystals generally provide excellent energy resolution and thus have been widely used in particle physics experiments for decades. Commonly used scintillators in calorimeters include NaI(Tl), pure CsI, CsI(Tl), BGO, PbWO4, and others [25]. Table 1 summarizes the properties of these inorganic scintillators. NaI(Tl) was widely used in the past decades due to its low cost and high light output. Nonetheless, it is hygroscopic and has a relatively low density, making it less suitable for large-scale particle physics experiments at present. Despite its short radiation length resulting in higher energy deposition efficiency, the low light output of BGO and PbWO4 leads to insufficient energy resolution. Although the pure CsI possesses a relatively fast decay time, its low light output leads to poor energy resolution. In the MACS experiment, the energy resolution of the pure CsI calorimeter for single $\gamma$ event and double $\gamma$ events was 69% and 52%, respectively [41]. Doping thallium (Tl) in pure CsI can significantly increase its light yield and result in excellent energy resolution. The MACE ECAL is required to maintain high precision for $e^{+}e^{-}$ annihilation $\gamma$-rays. Consequently, CsI(Tl) is chosen as the reference scintillator in this ECAL design.
Table 1: Properties of several inorganic scintillation crystals [42].
@@
  • $f$
    fast component
  • $s$
    slow component
The photosensor is also a crucial component in the calorimeter system. Its technical specifications (e.g., sensitive area, photon detection efficiency) determine the efficiency with which optical photons are converted into electrical signals that can ultimately be recorded. Commonly used photosensors in current particle and nuclear physics experiments include Photomultiplier Tubes (PMTs) and Silicon Photomultipliers (SiPMs) [43]. PMTs are relatively less expensive compared to SiPMs for the same sensitive area. Despite their limited quantum efficiency, high operating voltage, and shielding requirements in magnetic fields, PMTs are widely used in experiments. On the other hand, SiPMs offer excellent photon detection efficiency, a compact size, and tolerance to magnetic fields. However, they come with higher costs, dark current due to semiconductor thermal noise, and crosstalk between adjacent pixels [44]. Currently, PMTs are considered the benchmark photosensor in this paper.
@@ II.2 Optimization of Geometry Parameters
A detailed Monte-Carlo simulation of the module Type-HEX06, which is characterized by both the largest number and size, is performed using MACE offline software based on Geant4 v11.2.2. The physics processes are modeled with the reference physics list FTFP_BERT_EMZ, G4OpticalPhysics, and the Unified surface model. Additionally, the optical properties of Teflon-wrapped CsI(Tl) crystals (such as light yield, emission spectrum, decay time, etc.) and the PDE of the photosensor are considered based on manufacturer data sheets. Monoenergetic 511 keV $\gamma$-rays from a point-biased source are directed onto the center of a crystal optically coupled with a photosensor, which is registered as a Sensitive Detector (SD). The area of SD is adjusted to fully cover the outer surface of a Type-HEX06 crystal to estimate the intrinsic energy resolution of the crystal. The crystal length varies from 10 cm to 19 cm, and the inner radius of ECAL varies from 15 cm to 24 cm. Peak finding is performed using TSpectrum within the ROOT framework [46] to record the maximum count of scintillation photons spectra collected by the SD.
Meanwhile, the detection efficiency is given by the ratio of the number of events with energy greater than 500 keV to the total number of incident particles. The results of scanning each geometry parameter combination are detailed in Fig. 3.
@@
It should be noted that there is a trade-off between energy resolution and detection efficiency. The result demonstrates that increasing of the crystal volume causes a decrease in light collection efficiency and a deterioration in energy resolution, as investigated by Refs. [47, 48]. It is reasonable to assume that, as optical photons travel a longer distance, some of them will be absorbed by the material directly or after multiple reflections, resulting in reduced light collection efficiency. On the other hand, the crystal length has a limited effect on the detection efficiency because the longitudinal leakage of 511 keV $\gamma$-rays is negligible. Incident particles are more likely to escape laterally due to Compton scattering; thus, detection efficiency increases with the ECAL inner radius, which affects the module width. To balance the energy resolution and the detection efficiency, a Figure-of-Merit (F.o.M) function is constructed as follows:
@@
where $N_{pho}$ is the maximum count of scintillation photons spectra, $\eta$ is the detection efficiency. The parameters are ultimately chosen to be a crystal length of 10 cm and an ECAL inner radius of 20 cm based on consideration of both F.o.M and spatial constraints.
In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the ECAL chamber. Some modules must be removed to ensure that the incoming signal positrons are not obstructed. Specifically, the beam windows are set with an upstream radius of 50 mm and a downstream radius of 5 mm. Consequently, 19 modules at the beam entrance and 1 module at the exit are removed, resulting in a total of 622 modules and achieving 97% coverage of the $4\pi$ solid angle.
Figure 4: Calibration by Geant4 simulation of Type-HEX06 module. The red line is the FWHM curve fit by simulation data (blue dots), while the green line is the fit energy resolution.
The class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu. The crystal is coupled with a 1.5” diameter PMT. An isotropic point source, wrapped in 1 mm thick plastic, is placed at the entrance of the module. The energy dependence of FWHM is fit using  Eq. 2:
$\text{FWHM}=a+b\sqrt{E_{0}+cE_{0}^{2}}~{},$ (2)
where the best-fit value of $a$, $b$, $c$ are -1.56, 1.82, $1.72\times 10^{-4}$, respectively. The fit FWHM and energy resolution results of each full-energy peak are plotted in Fig. 4. For Type-HEX06 modules, the energy resolution is 9% at 511 keV, which is noticeably better than the resolution achieved by the calorimeter used in the MACS experiment [41]. Future improvements could include the adoption of large-area SiPM arrays with higher PDE.
@@ II.3.1 M-to-$\bar{\text{M}}$ Conversion Signal
  • Threshold cut: $E_{thr}>0.2~{}\text{MeV}$;
  • Energy difference cut: $\Delta E_{recon}<0.1~{}\text{MeV}$;
  • Angle cut: $\theta_{recon}>160^{\circ}$.
These criteria stand for trigger threshold, energy difference, and angle of two reconstructed hits, respectively. Finally, the ECAL response spectra are smeared using the FWHM curve described by Eq. 2. Events that fall within the $3\sigma$ interval of the full-energy peak are identified as signal events.
Figure 5: Spectra of original and reconstructed annihilation signals of M-to-$\bar{\text{M}}$ conversion by simulation.
Fig. 5 summarizes the simulated spectra of the positron annihilation signal. In the original $\gamma$ spectrum, the 511 keV full-energy peak accounts for only about 30% of the total counts. This is attributed to some of the $\gamma$-rays scattering to adjacent modules after depositing part of their energy, which forms a Compton continuum. Additionally, a peak at approximately 80 keV appears in the single $\gamma$ spectrum. This peak is attributed to characteristic X-rays emitted by lead atoms in the MCP. As a result, the signal efficiency is significantly reduced. However, by clustering the energy responses from each event, the total energy of the scattered photons can be reconstructed. As can be seen from Fig. 5, the energy has been reconstructed effectively. The expected signal is thus represented by the $3\sigma$ interval of 1.022 MeV.
Table 2: Contributions of signal Efficiency.
@@
Geometry 97%
No cut 55.5%
Energy difference cut 55.3%
Threshold cut 49.9%
Angle cut 43.5%
Total 42.8%
The result of the simulation indicates a 42.8%-55.5% signal efficiency in $3\sigma$ for ECAL. Losses of signal efficiency are as detailed in Table 2. Applications of various event selection criteria will depend on the real background level. The energy resolution and signal efficiency of ECAL are both superior to those of the MACS calorimeter, which is expected to significantly enhance the sensitivity of MACE.
II.3.2 Beam Background

### v3

@@ Abstract
The Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for charged lepton flavor violation and increase the sensitivity by more than two orders of magnitude compared to the MACS experiment at PSI in 1999. A clear signature of this conversion is the positron produced from antimuonium decay. This paper presents a near-$4\pi$-coverage calorimeter designed for MACE, which can provide an energy resolution of 10.8% at 511 keV, and a signal efficiency of 78.3% for annihilation $\gamma$-ray events. Detailed Monte-Carlo simulations using MACE offline software based on Geant4 are performed for geometry optimization, coincidence system design, background estimation, and benchmark detector validation.
@@ I Introduction
Muonium (M) is an exotic atom consisting of a bound state of $\mu^{+}e^{-}$, while its antiatom, antimuonium ($\bar{\text{M}}$), is composed of $\mu^{-}e^{+}$. From the perspective of the Standard Model Effective Field Theory (SMEFT), the spontaneous M-to-$\bar{\text{M}}$ conversion ($\mu^{+}e^{-}\rightarrow\mu^{-}e^{+}$), which violates lepton flavor by two units instead of one, can probe distinct effective operators and thereby constrain a different parameter space [8, 9, 10]. As a result, the key innovation in investigating M-to-$\bar{\text{M}}$ conversion is providing additional constraints on specific new physics models, distinct from other CLFV processes [11, 12, 13]. The most recent measured upper limit ($8.3\times 10^{-11}$ at 90% C.L.) for the conversion probability in a 0.1 T magnetic field was reported by the MACS experiment at PSI in 1999 [14, 15]. Another experiment focused on M-to-$\bar{\text{M}}$ conversion was proposed and conceptually designed at PRISM in 2003, as reported in [16]. There is also a proposal at J-PARC in Japan to search for this process using a novel scheme [17]. But none of these have progressed to further studies. Surprisingly, the limit of M-to-$\bar{\text{M}}$ conversion has remained unchallenged for two decades. There is an urgent need for a new-generation experiment.
In light of significant advancements in muon beamlines and particle detector techniques nowadays, the Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for M-to-$\bar{\text{M}}$ conversion and has the potential to improve the upper limit by three orders of magnitude [18, 19]. With the construction of the China advanced NUclear physics research Facility (CNUF) and other local muon sources [20, 21, 22, 23], MACE is expected to be the first muon-CLFV experiment in China, offering unique perspectives compared to other related experiments, including those at colliders. The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Positron Transport System (PTS), a Microchannel Plate (MCP) [24, 25], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence. The specifications of muon beamline in MACE is expected to produce a surface muon flux on target of $10^{8}~{}\mu^{+}/$s, with a repetition rate of 30–50 kHz and a pulse width duty cycle below 20%. A schematic of the MACE apparatus is shown in Fig. 1.
@@
In the MACE concept, surface muons with a momentum of 26.3 MeV$/c$ are stopped at a production target where M is formed and then possibly diffuses into the vacuum and converts into $\bar{\text{M}}$. The decay products of $\bar{\text{M}}$ include a Michel electron with an energy up to 52.8 MeV and a positron with atomic-shell energy ($\langle E_{\text{k}}\rangle=13.5$ eV). The Michel electron will be reconstructed by MMS which is expected to achieve $10^{-4}$ level suppression of background in the ECAL end through preliminary simulation. On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the PTS. The transverse position of the $e^{+}$ is conserved during the transportation to MCP, with a flight time FWHM of 6.9 ns. The cylindrical magnet with a field strength of 0.1 T outside ECAL is used to constrain the positron beam spot after transportation. While the MCP, with a timing resolution of 500 ps, is hit by a positron, the ECAL will be triggered and measure the produced two back-to-back annihilation $\gamma$-rays. As a rare-process searching experiment, MACE requires high sensitivity for the annihilation $\gamma$-rays to discriminate the signal from backgrounds and retain as much event information as possible. Thus, an ECAL with good energy resolution and high efficiency is essential for the design of MACE.
@@
The main purpose of this study is to propose a design for the MACE $4\pi$ calorimeter using the Monte-Carlo simulation method. In this work, the geometry of ECAL is generated and optimized using the MACE offline software based on the Geant4 toolkit [35, 36]. Subsequently, the energy resolution, signal efficiency, and background level of ECAL are estimated, providing a reference for the MACE conceptual design. This paper is structured as follows: Section I introduces the MACE calorimeter design and geometry optimization. Section II presents the simulation results for both signal and background. Section III describes the experiment setups of a benchmark detector, specifically the assembly used for an initial test of the CsI(Tl) crystal. The paper concludes with a discussion of the conclusions and perspectives for future work in Section IV.
@@ II.1 Overview
To meet the requirements of enhanced sensitivity (e.g., an excellent energy resolution, a high detection efficiency, and a good geometric acceptance), we propose a near-$4\pi$-coverage polyhedral calorimeter design. In this section, we present a design scheme for ECAL, including details on the selection of scintillation materials and photosensors. This proposed design will serve as a basis for the simulation work discussed in the following sections.
@@
Table 1: Properties of several inorganic scintillation crystals [41, 42]
@@
  • f fast component, s slow component
Calorimeters with inorganic scintillation crystals generally provide excellent energy resolution and thus have been widely used in particle physics experiments for decades. Commonly used scintillators in calorimeters include NaI(Tl), pure CsI, CsI(Tl), BGO, PbWO4, and others [43]. Table 1 summarizes the properties of these inorganic scintillators. The selection of scintillation crystal for MACE ECAL is based on the energy region of interest less than 1 MeV. NaI(Tl) was widely used in the past decades due to its low cost and high scintillation light output. Nonetheless, it is hygroscopic and has a relatively low density, making it less suitable for large-scale particle physics experiments at present. The problem of hygroscopy is identical for LaBr3(Ce). Moreover, considering the overall budget of this experiment, ultrafast and bright crystals such as LYSO(Ce), GAGG(Ce), or LaBr3(Ce) are not currently included but may be considered for potential future upgrades. The low light output of PbWO4 leads to insufficient energy resolution in the low-energy regime. The same is true for pure CsI. In the MACS experiment, the energy resolution of the pure CsI calorimeter for single $\gamma$ event and double $\gamma$ events was 69% and 52%, respectively [44, 14]. Doping thallium (Tl) in pure CsI can significantly increase its light yield, making it one of the brightest scintillators known. The objective of MACE ECAL is is to achieve excellent energy resolution for $\gamma$-rays produced in $e^{+}e^{-}$ annihilation. This requires optimizing the light yield of the scintillator, which is a key factor in precise energy measurements. Consequently, CsI(Tl) is chosen as the reference scintillator in this ECAL design.
The photosensor is also a crucial component in the calorimeter system. Its technical specifications (e.g., sensitive area, photon detection efficiency) determine the efficiency with which optical photons are converted into electrical signals that can ultimately be recorded. Commonly used photosensors in current particle and nuclear physics experiments include Photomultiplier Tubes (PMTs) and Silicon Photomultipliers (SiPMs) [45]. PMTs are relatively more cost-effective than SiPMs for the same sensitive area but are constrained by the low quantum efficiency (QE), high operating voltage, and shielding requirements in magnetic fields. On the other hand, SiPMs offer excellent photon detection efficiency (PDE), a compact size, and tolerance to magnetic fields. However, they come with higher costs, dark current due to semiconductor thermal noise, and crosstalk between adjacent pixels [46]. Currently, PMTs are considered the benchmark photosensor in this paper. A preliminary test of the SiPM performance is also discussed in Section III.2.
@@ II.2 Optimization of Geometry Parameters
A detailed Monte-Carlo simulation of the module Type-HEX06, which is characterized by both the largest number and size, is performed using MACE offline software based on Geant4 v11.2.2. The module mentioned above and below is defined as a combination of a scintillation crystal and a photosensor assembled together. The physics processes are modeled with the reference physics list QBBC_EMZ, G4OpticalPhysics, and the Unified surface model. Additionally, the optical properties of Teflon-wrapped CsI(Tl) crystals (such as light yield, emission spectrum, decay time, etc.) and the PDE of the photosensor are considered based on manufacturer data sheets. Detailed configurations of the simulation are presented in Table 2. Note that to better simulate the energy resolution of CsI(Tl), a Fano Factor $F_{\text{N}}$ is introduced [48, 49], and its squared value is assigned to the RESOLUTIONSCALE parameter in Geant4. Monoenergetic 511 keV $\gamma$-rays from a point-biased source are directed onto the center of a crystal optically coupled with a photosensor, which is registered as a Sensitive Detector (SD). The area of SD is adjusted to fully cover the outer surface of a Type-HEX06 crystal to estimate the intrinsic energy resolution of the crystal. The crystal length varies from 10 cm to 19 cm, and the inner radius of ECAL varies from 15 cm to 24 cm. Peak finding is performed using TSpectrum within the ROOT framework [50] to record the maximum count of scintillation photons spectra collected by the SD. Meanwhile, the detection efficiency is given by the ratio of the number of events with energy greater than 500 keV to the total number of incident particles. The results of scanning each geometry parameter combination are detailed in Fig. 3.
Table 2: Specification of material properties.
Material Density Refractive index Absorption length Reflectivity
$\rho$ (g$/\text{cm}^{3}$) $n$ $\lambda$ (cm) $R$
Crystal CsI(Tl) 4.51 1.79 37 -
Reflector Teflon 2.25 1.35 - 0.98
Coupler EJ-550 0.97 1.46 100 -
PMT window Glass 2.40 1.49 -
SiPM window Epoxy 1.18 1.55
PMT cathode Bialkali 2.00 - 0
SiPM sensor Silicon 2.33
@@
It should be noted that there is a trade-off between energy resolution and detection efficiency. The result demonstrates that increasing of the crystal volume causes a decrease in light collection efficiency and a deterioration in energy resolution, as investigated by Refs. [51, 52]. It is reasonable to assume that, as optical photons travel a longer distance, some of them will be absorbed by the material directly or after multiple reflections, resulting in reduced light collection efficiency. On the other hand, the crystal length has a limited effect on the detection efficiency because the longitudinal leakage of 511 keV $\gamma$-rays is negligible. Incident particles are more likely to escape laterally due to Compton scattering; thus, detection efficiency increases with the ECAL inner radius, which affects the module width. To balance the energy resolution and the detection efficiency, a Figure-of-Merit (F.o.M) function is constructed, following the function used in Ref. [53]:
@@
where $N_{\text{pho}}$ is the maximum count of scintillation photons spectra, $\eta$ is the detection efficiency. These two quantities are normalized by their mean values. The geometric parameters are ultimately chosen to be a crystal length of 10 cm and an ECAL inner radius of 20 cm based on consideration of both F.o.M and spatial constraints.
In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the region of positron detection system. Some modules must be removed to ensure that the incoming signal positrons are not obstructed. Specifically, the beam windows are set with an upstream radius of 50 mm and a downstream radius of 5 mm. Consequently, 19 modules at the beam entrance and 1 module at the exit are removed, resulting in a total of 622 modules and achieving 97% coverage of the $4\pi$ solid angle. For an estimation of energy resolution, the class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu. Instead of the fully-covered SD in pervious simulation, the crystal is coupled with an 1.5-inch diameter PMT volume. An isotropic point source, wrapped in 1 mm thick plastic, is placed at the entrance of the module. The energy dependence of FWHM is fit using Eq. 2:
$\text{FWHM}=a+b\sqrt{E_{0}}~{},$ (2)
where the best-fit value of $a$, $b$ are $-7.4707$ and $2.7638$, respectively. The fit FWHM and energy resolution results of each full-energy peak are plotted in Fig. 4. For the Type-HEX06 modules, the energy resolution is 10.8% at 511 keV, which is 6–7 times better than the resolution achieved by the calorimeter used in the MACS experiment.
Figure 4: Calibration by Geant4 simulation of Type-HEX06 module. The red line is the FWHM curve fit by simulation data (blue dots with error bars), while the green line is the fit energy resolution. The errors are introduced by fitting the full-energy peak.
@@ II.3.1 M-to-$\bar{\text{M}}$ conversion signal
  • Trigger threshold: $E_{\text{Thr}}>2\sigma$;
  • Energy difference cut: $\Delta E_{\text{Recon}}<3\sigma$;
  • Angle cut: $\theta_{\text{Recon}}>170^{\circ}$.
These criteria stand for trigger threshold, energy difference, and angle of two reconstructed tracks, respectively. Finally, the ECAL response spectra are smeared using the FWHM curve described by Eq. 2. Events that fall within the $3\sigma$ interval of the full-energy peak are identified as signal events.
Figure 5: Spectra of single Type-PEN module (blue line), single Type-HEX module (green line) and reconstructed annihilation signals (red line) of M-to-$\bar{\text{M}}$ conversion by simulation.
Fig. 5 summarizes the simulated spectra of the positron annihilation signal. In the original single-module $\gamma$ spectrum, the 511 keV full-energy peak accounts for only 47.2% of the total counts. This is attributed to some $\gamma$-rays scattering to adjacent modules after depositing part of their energy, which forms a Compton continuum. As a result, the signal efficiency will be significantly reduced. However, by clustering the energy responses from each event, the total energy of the scattered photons can be reconstructed. As can be seen from Fig. 5, the energy has been reconstructed effectively. The expected signal is thus represented by the $3\sigma$ interval of 1.022 MeV.
Table 3: Summary of factors contributing to the signal efficiency for double $\gamma$ events of the ECAL.
@@
Geometric efficiency of single $\gamma$ event $\varepsilon^{1\gamma}_{\text{Geom}}$ 95.2%
Geometric efficiency of double $\gamma$ events $\varepsilon^{2\gamma}_{\text{Geom}}$ 93.6%
Energy over threshold $\varepsilon_{\text{Trig}}$ ($E_{\text{Thr}}>2\sigma$) 98.7%
Reconstruction efficiency $\varepsilon_{\text{Recon}}$ $\Delta E_{\text{Recon}}<3\sigma$ 90.8%
$\theta_{\text{Recon}}>170^{\circ}$ 86.0%
Total signal efficiency 78.3%
The result of the simulation indicates a 78.3% signal efficiency for double $\gamma$ events in total. The factors contributing to the total signal efficiency are summarized in Table 3, where the standard deviation in the full-energy peak $\sigma$ is determined by the energy resolution. These results suggest that the efficiency performance of the ECAL surpasses that of the MACS calorimeter, which achieves an $\varepsilon^{1\gamma}$ of 79%.
II.3.2 Beam-related backgrounds